OpenStudy (anonymous):

The force of gravity on the moon is approximately one sixth that of the earth. The direct variation formula for weight on the earth compared to weight on the moon would be e = 6m, where e = weight on the earth and m = weight on the moon. What would be the weight of a 180 pound man on the moon?

OpenStudy (anonymous):

e = 6m, e = 180 m = e/6 = 180/6 = 30 pound